OpenStudy (anonymous):

Answer fast please :) Which of these ordered pairs is a solution to the inequality y – 2x ≤ –3? A. (2, 4) B. (1, –1) C. (3, 4) D. (–2, 3)

OpenStudy (anonymous):

Hey XD

OpenStudy (anonymous):

D

OpenStudy (anonymous):

heyyy lolXD

OpenStudy (anonymous):

ooh Pegasus ;D

OpenStudy (anonymous):

Can anyone help me with sample size?

OpenStudy (texaschic101):

its easy to check to see if it is D...just sub in your points in the inequality and see if it is true. y - 2x <= -3 (using points (-2,3) 3 - 2(-2) <= -3 3 + 4 <= -3 7 <= -3 (this is not true) D is not correct. But now that you know how to do it, you can find the correct answer.

OpenStudy (anonymous):

@texaschic101 may you please help me with finding sample size

OpenStudy (texaschic101):

I am not too good with sample size, but I can look at the problem

OpenStudy (anonymous):

Okay. It says to find the sample size required to achieve the given margin of error. 1. +- 7% 2. +- 6%

OpenStudy (texaschic101):

I am so sorry...I do not know the sample size stuff

OpenStudy (anonymous):

:(

OpenStudy (anonymous):

What about normal distribution?

OpenStudy (texaschic101):

y - 2x <= -3 --- using (2,4) 4 - 2(2) <= -3 4 - 4 <= -3 0 <= -3 correct....0 is LESS then or equal to -3 answer is (2,4)

OpenStudy (anonymous):

Here's a picture of this inequality if it helps you visualize it:

OpenStudy (anonymous):

Thanks :) I did 1,-1 and got it wrong BUT got an 80% thanks everyone :D

OpenStudy (anonymous):

You got it wrong??? How could you?? I'm kidding, glad you got an 80% grade. :)

OpenStudy (anonymous):

xD I thought when it said <= and it equals the same thing I would be good but I shouldve kept working it out